PINFO(1) ContributedSoftware PINFO(1) NAME pinfo−playInfocomInc.adventures SYNOPSIS pinfo[−aAehoOpstvV][−ccontext][−iindent][−llines][−mmargin][−rsavefile][displayflags][datafile[.dat]] DESCRIPTION ThisprogramallowsmostofthepopularInfocomtextadventurestobeplayedonyoursystem. ThegamesmustbeStandardSeriesfortheinterpretertofunctioncorrectly;seethefile INFOTBL.TXTintheproductreleaseforfulldetailsonallInfocomgames. The−hoptionwilltellyouwhatversionofZ-CodeInterpreteryouneedforthatadventure.The informationalprintingoptions(−h,−o,−O,−v,and−V)willworkwithanyadventure,butyou won’tbeabletoactuallyplayanygamesexceptforStandardSeriesadventures. OPTIONS Thecommand-lineargumentsallowforscreencustomizations,andalsovariousformsofcheating and/ordebugging. Informational Theseoptionwillprintinformationonthedatafiletothescreen.Ifanyofthesearegiventhe gameisnotactuallyplayed,informationisjustprinted. Notethatalltheseoptionsexcept−hand−Vwillprintinformationthatyouarenotsupposed toknow:examinationofthisinformationonagameyouhaven’tsolvedcouldseriouslydetract fromyourenjoymentofthegame. −hprintthedatafileheader −oprinttheobjectnames,attributesandlinks −Oprinttheobjecttree −vprintthegamevocabulary −Vverifythegamedatafiletoensureithasn’tbeencorrupted(togetthissamefeaturewhile playingusethe$verifycommand) GamePlay Theseoptionsaffectdifferentaspectsofgameplay.Someofthemmaydefinitelybeconsidered cheatingastheyprintinformationabouttheinternalsofthegameasyouareplayingit. −adisplaymodificationstoobjects’attributeswhileplaying −Adisplayvaluetestsofobjects’attributeswhileplaying −ccontext setthenumberoflinesofcontexttokeepatthetopofthescreenwhenpaginglongoutput(defaultis2) −eechoeachcommandbeforeexecuting −iindent setthenumberofspacesintheleftmargin(defaultis0).Notethatonlygameoutput textwillbeindented;inputpromptsandinputwillnot. −mmargin setthenumberofspacesintherightmargin(defaultis2) −pdon’tpagelongoutput −Psetthealternatepromptflag(thisflagisnotusedinallgames) Revision3.0 21October1992 1 PINFO(1) ContributedSoftware PINFO(1) −rsavefile causestheinterpretertorestorethesavedgamesavefileafterstartingup. −sdon’tprintstatusline −tdisplaymodificationsoftheobjecttreewhileplaying −TsettheTandyLicensingflag(thisflagisnotusedinallgames).Onsomegamesthis causesprintedtexttobeslightlydifferent. InterpreterEscape Someoftheaboveflagsandoptionsmayalsobemodifiedduringgameplayviainterpreter escapes.Anycommandbeginningwiththeescapecharacter(normally‘‘@’’)isassumedtobean interpreterescapecommand.Thecurrentcommandsandtheirvaluesmaybelistedbygivingjust theescapecharacterbyitself.Currentcommandsinclude: attrTogglestheattributeprintingvariable(see−aabove) echoTogglesinputlineechoing(see−eabove) pagerTogglespagedoutputmode(see−pabove) prompt Togglesthealternateprompt(see−Pabove) statusTogglesstatuslineprinting(see−sabove) tandyTogglestheTandylicenseflag(see−Tabove) xfersTogglestheobjecttransferflag(see−tabove) Additionalescapecommandsmaybedefinedbyindividualterminalinterfaces;seebelow. TERMCAP Thissectiongivesinformationspecifictothetermcapterminalinterface.Thisinterfaceisforuse withUNIXmachinesorotherswithUNIX-compatibletermcaporterminfolibraries.Inaddition totermcapandterminfoitalsosupportsthetermio,termios,andsgttyline-disciplinelibraries. Options Theseoptionsareinadditiontotheabove,generaloptions: −Cfile Specifiesafileofcompletionwords.Ifnotgivennowordsareavailableatstartupand anywordsdefinedduringplaywillnotbestored.Notepinfomustbecompiledwith GNUReadlinesupporttousethisoption. −Hfile specifiesaGNUReadlinelibraryhistoryfiletobeused.Ifnotgiven,historyisnotstored inafile.NotepinfomustbecompiledwithGNUReadlinesupporttousethisoption. Subshells TheTERMCAPversionoftheinterpretercaninvokecommandsinasubshelloraninteractive subshellinthenormalUNIXfashion:i.e.,ifacommand’sfirstcharacterisanexclamationmark (bang)(‘‘!’’)thentheremainderofthecommandispassedverbatimtoasubshell.Ifthereisno textaftertheexclamationmarkthenaninteractivesubshellisinvoked. ThevalueoftheenvironmentvariableSHELLisusedasthesubshell,or/bin/shifSHELL doesn’texistorisempty. SeetheGNUReadlinesectionforanexceptionwhenusingGNUReadline. GNUReadline TheTERMCAPversionoftheinterpretercontainsoptionalsupportfortheGNUReadline library.Thislibraryallowsinputlineeditingandhistoryretrieval,includingfullincremental searchingandcsh-stylecommand-lineediting.BydefaultthekeybindingsarethoseofGNU Emacs,butthelibraryisfullyflexibleandcanbeconfiguredtoworkinvimode,oryoucanroll-Revision3.0 21October1992 2 PINFO(1) ContributedSoftware PINFO(1) your-ownbindings. Becausecsh-styleeditingusesanexclamationmarkasthefirstcharacterforhistorysubstitutions, inordertogetasubshellcommand(seeabove)youshouldescapetheexclamationmarkwitha backslash(‘‘\’’)whenusingtheGNUReadlinelibrary.Forexample: >\!ls-l InadditiontothestandardGNUReadlinebindings,thesinglecharactercommand?willcause theinterpretertoprintthecurrenthistorylisttothescreen. SeetheGNUReadlinemanualforfulldetails. Completion IftheGNUReadlinelibraryisbeingused,thencompletionisavailablebypressingtheTABkey. AlistofallpossiblecompletionsisavailablebypressingTABtwiceinarow. Duringnormalgameplaytheinterpretermaintainsalistofpossiblewordstocompleteon.This listisreadinfromafilespecifiedwith−C(above)atstartupandwrittenouttothatfileatthe endofthegame.Modificationstothelistarepossibleviaspecialinterpreterescapecommands (seebelow). Filenamecompletionisavailableduringsave/restoreprocessing,whenevertheinterpreterisasking forafilename. InterpreterEscapes Twonewescapecommandsaredefinedinadditiontothenormalescapesifpinfoiscompiledto usetheGNUReadlinelibrary.Theseallowadditionanddeletionofwordsinthecommandcompletionlist. add<word> Add<word>tothecompletionlist.Duplicatesarenotallowed. del<word> Delete<word>fromthecompletionlist. EnvironmentVariables WhenusingtheinterpreterwithTERMCAPsupport,thefollowingenvironmentvariablesare used: SHELL Ifset,containsthefullpathoftheusersshell.Thisprogramisinvokediftheshellcommandistypedatthegamepromptwithnoarguments. TERM Thisvariablecontainstheterminaltypeusedtodisplaytheoutput. PIFLAGS Thisvariableshouldcontainflagswhichyouwouldalwayslikepresentonthepinfocommandline.Anyflagsgivenherewillbeoverriddenbyflagsactuallyonthecommandline, andbethenexttwovariablesinthecaseof−Hand−C. PIHIST ContainsthefilenameofthepinfohistoryfileusedwithGNUReadlinesupport(see above).Setting−Honthecommandlineoverridesthevalueofthisvariable. PIWORD Containsthefilenameofthepinfocommand-completionwordfileusedwithGNURead-linesupport(seeabove).Setting−Conthecommandlineoverridesthevalueofthis variable. STREAM Therearenostreamterminalinterfacespecificflagsorcapabilitites. Revision3.0 21October1992 3 PINFO(1) ContributedSoftware PINFO(1) DATAFILES Infocomgamesconsistofadatafileandaninterpreterwhichexecutestheinstructionscontained inthedatafiletorunthegame.Thisprogramisessentiallyaportabledatafileinterpreter. Ifyouhavethe‘‘LostTreasuresofInfocom’’packagefortheIBMorAmigayoushouldhavethe datafilesreadilyavailableonthedisks.IfyouhaveanyInfocomgamefortheMacintoshthen whentransferringtheprogramtoyourUNIXhostbesuretonotsendtheMacintoshresource sections(i.e.,don’tenabletheMacBinaryoption).MacInfocomprogramssanstheirresources arethedatafilesneededforusewiththisprogram. BUGS Someofthedatafileshaveerrorsintheirvocabularylists.Ifthe−voptiondetectssuchanerror itwillputsquarebrackets([])aroundtheoffendingword.Suchawordmaynotbeusedinthe game,sincetheinterpreterwillnotrecognizeit.Itactuallycan’tbeusedwith*any*interpreter, eventheonesfromInfocomitself;thesearebugsinthedatafile. AUTHORS TheoriginalsourcecodeforthisinterpreterwastheInfoTaskForce interpreter,versionREVC. PaulD.Smith<paulsmith@dg.com> ModifiedfromITFREVCtothecurrentversion.Manyenhancements:seeNEWSin thesourcedistributionforacompletelistofuser-visiblemodifications. OlafBarthel<olsen@sourcery.mxm.sub.org> ContributedtheAmigaterminalinterface. MikeThreepoint<linhart@remus.rutgers.edu> ContributedtheMS-DOSterminalinterface. COPYRIGHTS ThispinfointerpreterisdistributedundertheGNUGeneralPublicLicense(seethefileCOPY-INGinthesourcedirectory). Infocom,Inc.hasgoneoutofbusiness,buttherightstotheirgameshavebeensold(various times!)andarecurrentlystillowned.AllInfocomgamesarestillprotectedandassuchmaynot becopied. SomegamesreleasedbyInfocomhaveaveryrestrictivecopyrightwhichstatesthatthegamemay notbeplayedonanyplatformexceptthatforwhichitwassold.Pleaserefertothecopyright noticecontainedinyourgametodecidewhetheryouwillbeviolatingcopyrightlawsbeforeusing itwiththisinterpreter. NowthatLostTreasuresisavailablethereisnoexcusefornotpurchasingtheInfocomgames yourself;atabout$2.50to$3.00agame,includingtax,shipping,handling,manual,maps,and hints[inLTOIIonly]thisisabargainthelikesofwhichyouprobablywon’tseeagaininyour lifetime!:-). Let’sallsupportMediagenicintheirreleaseofLostTreasuresI&IIandmaybethey’llreleasethe restofthelostInfocomgames. Revision3.0 21October1992 4